# Data-light Uncertainty Set Merging with Admissibility: Synthetics, Aggregation, and Test Inversion

### v2

@@ Abstract
This article introduces a Synthetics, Aggregation, and Test inversion (SAT) approach for merging diverse and potentially dependent uncertainty sets into a single unified set. The procedure is data-light, relying only on initial sets and control levels, and it adapts to any user-specified initial uncertainty sets, accommodating potentially varying coverage levels. SAT is motivated by the challenge of integrating uncertainty sets when only the initial sets and their control levels are available—for example, when merging confidence sets from distributed sites under communication constraints or combining conformal prediction sets generated by different algorithms or data splits. To address this, SAT constructs and aggregates novel synthetic test statistics, and then derive merged sets through test inversion. Our method leverages the duality between set estimation and hypothesis testing, ensuring reliable coverage in dependent scenarios. A key theoretical contribution is a rigorous analysis of SAT’s properties, including a proof of its admissibility in the context of deterministic set merging. Both theoretical analyses and empirical results confirm the method’s finite-sample coverage validity and desirable set sizes.
@@ 1 Introduction
Uncertainty sets, such as confidence intervals and prediction intervals, are pivotal in statistical inference as they facilitate accurate representation and management of data variability. The integration of these sets has wide-ranging applications across various fields; however, considerable challenges emerge, especially when only the initial uncertainty sets and their control levels are available, along with possible intrinsic dependencies among the sets. To highlight the significance of set merging, we will first explore two prominent examples.
  1. (a).
@@
    Distributed Learning with Communication Constraint. In distributed learning, the primary objective is to collaboratively make inferences using data distributed across different studies. Recent advancements have focused on distributed mean estimation (Cai and Wei,, 2024), prediction (Humbert et al.,, 2023), and causal inference (Xiong et al.,, 2023). A major challenge in distributed learning is the presence of communication constraints, which can arise from bandwidth limitations, privacy concerns, or cost considerations. These constraints restrict the amount of information that can be exchanged between studies. In certain scenarios, local sites can only transmit the confidence set and its associated confidence level to a central aggregator, highlighting the necessity for effective and data-light methods to merge these uncertainty sets for robust and reliable inferences.
  2. (b).
@@
    Algorithmic Stability and Derandomization. Conformal prediction, pioneered by Vovk et al., (2005), has gained considerable popularity due to its minimal assumption requirements and its capability to provide finite-sample valid prediction sets for any black-box models. One of the most widely adopted versions is split conformal inference (e.g., Lei et al.,, 2018; Angelopoulos and Bates,, 2021), which is valued for its computational efficiency. However, the resulting prediction set can be influenced by the way the data is split. Additionally, variations in the prediction set may occur depending on the algorithm employed to compute the non-conformity score. To mitigate these issues, combining different prediction sets becomes a natural and necessary strategy.
@@
This paper aims to develop an efficient and flexible method to combine $L$ different–potentially dependent–uncertainty sets into a single set, given only the initial uncertainty sets and their corresponding control levels. We term this method a data-light approach, emphasizing that it requires no access to raw data and the processes used to construct the initial uncertainty sets. Formally, let $Y$ denote the prediction target, and let $\mathcal{C}_{\ell,\alpha_{\ell}}$ represent any initial uncertainty set from the $\ell$-th study such that
$\mathbb{P}(Y\notin\mathcal{C}_{\ell,\alpha_{\ell}})\leq\alpha_{\ell},\quad\ell=1,\ldots,L,$ (1)
@@
where $\alpha_{\ell}$’s are possibly varying control levels. The goal is to construct a merged set $\bar{\mathcal{C}}_{\alpha}$ such that $\mathbb{P}(Y\notin\bar{\mathcal{C}}_{\alpha})\leq\alpha$ for a pre-specified $\alpha\in(0,1)$, using only the available $\{(\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\}_{\ell=1}^{L}$, while also ensuring that the merged set remains small in size.
1.1 Related Work
@@ 1.2 Our Method and Contributions
In this paper, we develop a novel data-light procedure for merging uncertainty sets, which we term Synthetics, Aggregation, and Test inversion (SAT). The SAT procedure operates in three core steps. First, we propose a novel class of synthetic test statistics that depend solely on the initial uncertainty sets and their associated coverage levels. These statistics are designed to mimic the true, underlying unknown statistics used to construct the initial sets, crucially without requiring access to the original data. Second, these synthetic test statistics are aggregated from multiple input sets based on appropriate aggregation functions. Finally, the merged uncertainty set is derived through the test inversion of the aggregated synthetic statistics.
This data-light approach requires minimal assumptions and only uses the initial uncertainty sets and their coverage levels, making it broadly applicable across various scientific domains. Our work offers several key contributions:
  1. (a).
    We propose a principled framework for merging uncertainty sets based on the duality between hypothesis testing and set estimation. This allows us to convert the set merging problem into the aggregation of test statistics, providing robustness to arbitrarily dependent input sets.
  2. (b).
    We introduce the concept of “Synthetic Statistics”, which bypasses the need for raw data by effectively mimicking unknown oracle test statistics using only the available initial sets and their coverage levels.
  3. (c).
    Importantly, by proposing the idea of synthetic statistics, we provide a theoretical analysis that proves the admissibility of the SAT procedure in a specific context, from which the admissibility of the majority voting method, as a special case of SAT, is derived. This represents a significant theoretical contribution, providing foundational support for a commonly used heuristic.
  4. (d).
    We establish a finite-sample coverage probability guarantee for the merged set derived from the SAT procedure, valid without model assumptions, and analyze the asymptotic size properties of the merged set.
1.3 Organization
The rest of the paper is organized as follows. Section 2 describes the proposed SAT procedure in detail. In Section 3, we establish the coverage guarantee and admissibility of SAT and analyze the size of the merged set theoretically. Sections 4 and 5 explore the numerical performance through simulations and the ImageNet__\__val dataset. We conclude the paper with a discussion of future directions in Section 6. Further extensions of the methods and theories, along with technical proofs and additional numerical results, can be found in the Supplementary Material.
1.4 Notations
Denote by $[n]=\{1,\dots,n\}$ for $n\in\mathbb{Z}^{+}$, and let $\mathbb{R}^{\geq 0}$ be the set of non-negative real numbers. Let $\operatorname{\mathds{1}}(\cdot)$ denote the indicator function. Let ${\bm{1}}_{L}$ be an $L$-dimensional vector of 1’s. For two positive sequences $\{a_{n}\}_{n\geq 1},\{b_{n}\}_{n\geq 1}$, write $a_{n}=O(b_{n})$ if there exists a constant $C>0$ such that $a_{n}/b_{n}\leq C$ for all $n$. Let $\|f\|_{L_{2}([0,1])}=\left\{\int_{0}^{1}f(x)^{2}{\rm d}x\right\}^{1/2}$ be the $L_{2}$-norm of $f$ over $[0,1]$, and denote $\|f\|_{L_{1}([0,1])}=\int_{0}^{1}|f(x)|{\rm d}x$ as the $L_{1}$-norm of $f$ over $[0,1]$. Denote by $\Phi(\cdot)$ and $z_{1-\alpha}$ the cumulative distribution function (CDF) and the $1-\alpha$ quantile of a standard normal random variable, respectively.
2 The SAT Procedure
This section proposes a SAT approach for merging uncertainty sets, utilizing only the initial sets and their corresponding control levels, while ensuring a guaranteed coverage probability. A broader class of uncertainty sets, where the notion of error extends beyond the miscoverage rate, will be further discussed in Section C.1 of the supplement.
Inspired by the duality between hypothesis testing and set estimation (Casella and Berger,, 2024), our proposed SAT procedure involves the following three steps:
Step 1 (Synthetics): Derive synthetic test statistics from initial uncertainty sets.
Step 2 (Aggregation): Aggregate test statistics from different studies.
Step 3 (Test Inversion): Merge uncertainty sets via test inversion of aggregated test statistics.
This sequence of steps outlines the intuition behind the SAT procedure, as detailed in Algorithm 1. Each step will be explained further in the subsequent subsections.
Algorithm 1 SAT Procedure
@@ 2 The SAT Procedure
0:  The pairs $\{(\mathcal{C}_{\ell,\alpha_{\ell}},\alpha_{\ell})\}_{\ell\in[L]}$, candidate space $\mathcal{Y}$, a suitable aggregation function $G_{p}(\cdot)$ (or $G_{e}(\cdot)$), a control level $\alpha\in(0,1)$, an adjustment factor $\tau\in(0,1]$ (optional, $\tau$=1 by default).
0:  $\bar{\mathcal{C}}_{\alpha}\leftarrow\{\}$.
@@
3:        Generate synthetic p-values $p_{\ell}(y)$ by (2) (or synthetic e-values $e_{\ell}(y)$ by (3)).
@@
5:     Calculate $\bar{p}(y)=G_{p}\{\mathbf{p}(y)\}$ where $\mathbf{p}(y)=\{p_{1}(y),\ldots,p_{L}(y)\}$
                (or $\bar{e}(y)=G_{e}\{\mathbf{e}(y)\}$ where $\mathbf{e}(y)=\{e_{1}(y),\ldots,e_{L}(y)\}$).
6:     Update $\bar{\mathcal{C}}_{\alpha}\leftarrow\bar{\mathcal{C}}_{\alpha}\cup\{y\}$ if $\bar{p}(y)>\alpha$ (or if $\bar{e}(y)<\tau/\alpha$).
@@
2.1 Synthetic Statistics
A key component of the SAT procedure is the innovative construction of synthetic statistics, which translates initial uncertainty sets into test statistics through the duality of testing and interval estimation. Though the original data is not accessible, we can approximate the underlying (unknown) true statistics based on the initial sets and their levels, and construct synthetic versions of both p-values and e-values. In this section, we will detail this construction and examine the associated theoretical properties.
2.1.1 Synthetic p-value
A p-value $p\in[0,1]$ is a random variable that satisfies $\mathbb{P}(p\leq t)\leq t$ for all $t\in(0,1)$ under the null hypothesis [^2] With a slight abuse of terminology, the terms “p-values” and “e-values” below refer to both the random variables and their realized values, with the context clarifying the intended meaning.. It is well established that p-values can be employed to construct the uncertainty set, as formally stated in Propositions 7 and 8 in the subsequent sections. However, when only the uncertainty sets are provided, the true underlying statistics used to construct these sets, such as the p-values—referred to as “oracle p-values”—are not accessible. For each uncertainty set $\mathcal{C}_{\ell,\alpha_{\ell}}$ with control level $\alpha_{\ell}$, $\ell\in[L]$, we propose generating a random synthetic p-value that mimics such “oracle p-value” as follows:

### v3

@@ Abstract
This article introduces a Synthetics, Aggregation, and Test inversion (SAT) approach for merging diverse and potentially dependent uncertainty sets into a single unified set. The procedure is data-light, relying only on initial sets and their nominal levels, and it flexibly adapts to user-specified input sets with possibly varying coverage guarantees. SAT is motivated by the challenge of integrating uncertainty sets when only the initial sets and their control levels are available—for example, when merging confidence sets from distributed sites under communication constraints or combining conformal prediction sets generated by different algorithms or data splits. To address this, SAT constructs and aggregates novel synthetic test statistics, and then derive merged sets through test inversion. Our method leverages the duality between set estimation and hypothesis testing, ensuring reliable coverage in dependent scenarios. A key theoretical contribution is a rigorous analysis of SAT’s properties, including its admissibility in the context of deterministic set merging. Both theoretical analyses and empirical results confirm the method’s finite-sample coverage validity and desirable set sizes.
@@ 1 Introduction
Uncertainty sets, such as confidence intervals and prediction intervals, are pivotal in statistical inference as they facilitate accurate representation and management of data variability. The integration of these sets has wide-ranging applications across various fields. However, considerable challenges emerge, especially when only the initial uncertainty sets and their control levels are available, along with possible intrinsic dependencies among the sets. To highlight the significance of set merging, we will first explore two prominent examples.
@@
Example 1 (Distributed Learning with Communication Constraint). In distributed learning, the primary objective is to collaboratively make inferences using data distributed across different studies. Recent advancements have focused on distributed mean estimation (Cai and Wei,, 2024), prediction (Humbert et al.,, 2023), and causal inference (Xiong et al.,, 2023). A major challenge in distributed learning is the presence of communication constraints, which can arise from bandwidth limitations, privacy concerns, or cost considerations. These constraints restrict the amount of information that can be exchanged between studies. In certain scenarios, local sites can only transmit the confidence set and its associated confidence level to a central aggregator, highlighting the necessity for effective and data-light methods to merge these uncertainty sets for robust and reliable inferences.
@@
Example 2 (Algorithmic Stability and Derandomization). Conformal prediction, pioneered by Vovk et al., (2005), has gained considerable popularity due to its minimal assumption requirements and its capability to provide finite-sample valid prediction sets for any black-box models. One of the most widely adopted versions is split conformal inference (e.g., Lei et al.,, 2018; Angelopoulos and Bates,, 2021), which is valued for its computational efficiency. However, the resulting prediction set can be influenced by the way the data is split. Additionally, variations in the prediction set may occur depending on the algorithm employed to compute the non-conformity score. To mitigate these issues, combining different prediction sets becomes a natural and necessary strategy.
@@
This paper aims to develop an efficient and flexible method to combine $L$ different–potentially dependent–uncertainty sets into a single set, given only the initial uncertainty sets and their corresponding control levels. We term this method a data-light approach, emphasizing that it requires no access to raw data and the processes used to construct the initial uncertainty sets. Formally, let $Y$ denote the prediction target, and let $\mathcal{C}_{\ell}$ represent any initial uncertainty set from the $\ell$-th study such that
$\mathbb{P}(Y\notin\mathcal{C}_{\ell})\leq\alpha_{\ell},\quad\ell=1,\ldots,L,$ (1)
@@
where $\alpha_{\ell}$’s are possibly varying control levels. The goal is to construct a merged set $\bar{\mathcal{C}}_{\alpha}$ such that $\mathbb{P}(Y\notin\bar{\mathcal{C}}_{\alpha})\leq\alpha$ for a pre-specified $\alpha\in(0,1)$, using only the available $\{(\mathcal{C}_{\ell},\alpha_{\ell})\}_{\ell=1}^{L}$, while also ensuring that the merged set remains small in size.
@@
In this paper, we develop a novel data-light procedure which we term Synthetics, Aggregation, and Test inversion (SAT). The core of the SAT framework is to convert the problem of merging sets into a more tractable statistical problem of aggregating evidence. This is achieved in three steps. First, we construct novel synthetic test statistics, such as e-values or p-values, that depend solely on the initial uncertainty sets and their coverage levels. These statistics are designed to mimic the true, underlying unknown statistics used to construct the initial sets. Second, these synthetic test statistics are merged from the multiple input sets. Finally, the merged uncertainty set is derived through the test inversion of the aggregated synthetic statistic. A schematic overview of the SAT procedure is given in Figure 1.
[…]
Figure 1: A schematic illustration of the SAT procedure using synthetic e-values. For each candidate $y_{i}$, synthetic e-values are generated from the initial sets and are then aggregated. The final merged set, $\bar{\mathcal{C}}_{\alpha}$, is constructed by including only those candidates whose aggregated statistic passes a pre-defined significance threshold. In this illustration, the aggregated statistics for $y_{1}$ and $y_{m}$ pass the threshold (shaded green), while the one for $y_{2}$ does not (shaded red).
Our work makes several key contributions. First, we propose a principled framework that leverages the duality between hypothesis testing and set estimation to convert the set merging problem into one of aggregating test statistics, an approach that is robust to arbitrary dependencies among the input sets. Central to our framework is the concept of “synthetic statistics”, which mimic unknown oracle statistics while using only the initial sets and their coverage levels, thus bypassing any need for raw data. Second, the resulting merged set has a finite-sample theoretical coverage guarantee that is valid without any modeling assumptions. Finally, we establish SAT’s admissibility in the deterministic setting. This theoretical result provides foundational support for the commonly used voting heuristics. We establish not only that majority voting is a special case of our framework, but also that, more generally, any admissible procedure is equivalent to a weighted voting scheme. Our analysis therefore provides the first formal optimality guarantee for these popular methods, grounding a widely used family of heuristics in statistical admissibility theory.
Throughout the paper, we let $n]=\{1,\dots,n\}$ for an integer $n\in\mathbb{Z}^{+}$, and let $\mathbb{R}^{\geq 0}$ be the set of non-negative real numbers. The indicator function is denoted by $\operatorname{\mathds{1}}(\cdot)$, and $\mathbf{1}_{L}$ is an $L$-dimensional vector of ones. For two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ if the ratio $a_{n}/b_{n}$ is bounded. The $L_{1}$ and $L_{2}$ norms of a function $f$ over the unit interval are denoted by $\|f\|_{L_{1}([0,1])}=\int_{0}^{1}|f(x)|{\rm d}x$ and $\|f\|_{L_{2}([0,1])}=\{\int_{0}^{1}f(x)^{2}{\rm d}x\}^{1/2}$, respectively.
2 Deterministic Set Merging with Synthetic e-values
2.1 Synthetic e-value
We start with deterministic set merging procedures using synthetic e-values in this section, and will move to randomized processes with synthetic p-values in the following section.
Our procedure is motivated by the duality between set estimation and hypothesis testing, where a confidence set is formed by the collection of parameter values for which a null hypothesis is not rejected (Casella and Berger,, 2024). This principle allows us to reframe the problem of merging sets as one of aggregating test statistics.
Recently, the e-value has gained popularity as a measure of evidence for hypothesis testing (Vovk and Wang,, 2021; Shafer,, 2021). An e-value is a non-negative random variable whose expectation is at most one under the null hypothesis. A key property is that for any $\alpha\in(0,1)$, the test that rejects the null hypothesis if and only if $e\geq 1/\alpha$ controls the Type I error at level $\alpha$. This testing rule provides a natural way to construct a synthetic e-value from an existing uncertainty set. While the true “oracle e-value” used to form the initial set is unavailable in our data-light setting, we have the pair $(\mathcal{C}_{\ell},\alpha_{\ell})$ itself. The set $\mathcal{C}_{\ell}$ is, by construction, the non-rejection region of its underlying test. To align with the e-value framework, which rejects a hypothesis when $e\geq 1/\alpha_{\ell}$, we construct a statistic that emulates this behavior. Specifically, we assign a value of $1/\alpha_{\ell}$ to any point $y$ in the rejection region, i.e., $y\notin\mathcal{C}_{\ell}$, and zero otherwise. This directly motivates defining the synthetic e-value as:
$e_{\ell}(y)=\alpha_{\ell}^{-1}\cdot\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell}),\quad\forall y\in\mathcal{Y}.$ (2)
The function $e_{\ell}(\cdot)$ is referred to as an e-function. The validity of this construction is established in the following proposition.
Proposition 1.
2.2 Aggregating Synthetic e-values and Test Inversion
After constructing synthetic e-values for each initial set, the next step is to aggregate them into a single statistic, $\bar{e}(y)$. The appropriate aggregation method depends on the dependence structure of the initial sets.
We first consider the case where the synthetic e-values are mutually independent. This assumption holds, for instance, when $Y$ is a fixed parameter and each study independently collects data. A general aggregator for a vector $\mathbf{e}=(e_{1},\ldots,e_{L})$ of independent e-values is the average of products (Vovk and Wang,, 2021),
$\bar{e}_{k}=\binom{L}{k}^{-1}\sum_{\mathcal{I}_{k}\in\mathcal{B}_{k}}\prod_{\ell\in\mathcal{I}_{k}}e_{\ell},$ (3)
where $\mathcal{B}_{k}$ is the set of all $k$-element subsets of $[L]$. When $k$ is fixed, we abbreviate $\bar{e}_{k}$ to $\bar{e}$. As shown in the following proposition, if the initial e-values are valid, so is their aggregation.
Proposition 2.
Suppose $\mathbb{E}(e_{\ell}(Y))\leq 1$ for all $\ell\in[L]$. Let $\bar{e}_{k}(Y):=G_{e}\{\mathbf{e}(Y);k\}$ be defined as in (3). If the entries of $\mathbf{e}(Y)$ are mutually independent, then $\mathbb{E}\{\bar{e}_{k}(Y)\}\leq 1$ for any pre-determined $k\in[L]$.
If $Y$ is random, however, the synthetic e-values $e_{\ell}(Y)$ are generally not independent, as they share a common source of randomness. For this case of arbitrary dependence, a valid aggregation method is a convex combination.
Proposition 3.
Suppose $\mathbb{E}(e_{\ell}(Y))\leq 1$ for all $\ell\in[L]$. Let $\bar{e}(Y):=G_{e}\{\mathbf{e}(Y);\lambda\}=\sum_{\ell=1}^{L}\lambda_{\ell}e_{\ell}(Y)$, where $\lambda_{\ell}\geq 0$ for all $\ell$ and $\sum_{\ell=1}^{L}\lambda_{\ell}=1$. Then, we have $\mathbb{E}\{\bar{e}(Y)\}\leq 1$.
The final step is to transform the aggregated e-value $\bar{e}(y)$ back into a merged uncertainty set. This is achieved through test inversion (Casella and Berger,, 2024). An e-value $e(Y)$ controls the Type I error at level $\alpha$ using the test that rejects when $e(Y)\geq 1/\alpha$. Inverting this rule yields a $(1-\alpha)$-level uncertainty set. The following proposition formalizes this for our setting.
Proposition 4.
By combining Propositions 1 - 4, we see that for any pre-specified level $\alpha$, the final merged set can be constructed as
$\bar{\mathcal{C}}_{\alpha}=\{y\in\mathcal{Y}:\bar{e}(y)<\tau/\alpha\},$
where $\bar{e}(y)$ is a valid aggregated e-value. The complete Synthetics, Aggregation, and Test inversion (SAT) procedure is summarized in Algorithm 1. It provides a unified framework applicable to both the deterministic approach using e-values, as discussed, and a randomized version using synthetic p-values, which we introduce in Section 3. The validity of SAT is formalized in the following theorem.
Algorithm 1 The SAT Procedure
@@ 2.2 Aggregating Synthetic e-values and Test Inversion
0: The pairs $\{(\mathcal{C}_{\ell},\alpha_{\ell})\}_{\ell\in[L]}$, candidate space $\mathcal{Y}$, a suitable aggregation function $G_{e}(\cdot)$ (or $G_{p}(\cdot)$), a control level $\alpha\in(0,1)$, an adjustment factor $\tau\in(0,1]$ (optional, $\tau$=1 by default).
[…]
@@
3:   Generate synthetic e-values $e_{\ell}(y)$ by (2) (or synthetic p-values $p_{\ell}(y)$ by (4)).
@@
5:  Calculate $\bar{e}(y)=G_{e}\{\mathbf{e}(y)\}$ where $\mathbf{e}(y)=\{e_{1}(y),\ldots,e_{L}(y)\}$
      (or $\bar{p}(y)=G_{p}\{\mathbf{p}(y)\}$ where $\mathbf{p}(y)=\{p_{1}(y),\ldots,p_{L}(y)\}$).
6:  Update $\bar{\mathcal{C}}_{\alpha}\leftarrow\bar{\mathcal{C}}_{\alpha}\cup\{y\}$, if $\bar{e}(y)<\tau/\alpha$ (or if $\bar{p}(y)>\alpha$).
@@
Theorem 1.
Suppose the initial uncertainty sets $\{\mathcal{C}_{\ell}\}_{\ell=1}^{L}$ satisfy (1). Let the synthetic statistics be generated by (2) or (4), and let the aggregation function be one that produces a valid aggregated e-value or p-value, as established in Propositions 23 and 67. Then the merged set $\bar{\mathcal{C}}_{\alpha}$ produced by Algorithm 1, with the default choice of $\tau=1$, is a valid $(1-\alpha)$-level uncertainty set satisfying $\mathbb{P}(Y\in\bar{\mathcal{C}}_{\alpha})\geq 1-\alpha$ for any $\alpha\in(0,1)$.
Remark 1.
The SAT procedure with synthetic e-values and a convex combination is equivalent to a weighted voting scheme. A point $y$ is included in the merged set if $\sum_{\ell=1}^{L}\lambda_{\ell}e_{\ell}(y)<\tau/\alpha$. Substituting the definition of synthetic e-values, this condition becomes $\sum_{\ell=1}^{L}\frac{\lambda_{\ell}}{\alpha_{\ell}}\{\operatorname{\mathds{1}}(y\notin\mathcal{C}_{\ell})\}<\frac{\tau}{\alpha}.$ Here, each set $\mathcal{C}_{\ell}$ casts a “vote” for including $y$, and the vote is weighted by $\lambda_{\ell}/\alpha_{\ell}$. When $\alpha_{\ell}=\alpha/2$ for all $l\in[L]$ and an arithmetic mean is used for aggregation ($\lambda_{\ell}=1/L$), the majority voting procedurse of Gasparin and Ramdas, (2024) appears as a special case.
2.3 Admissibility of SAT for Deterministic Set Merging
In the previous subsection, we formally introduced the SAT procedure for deterministic set merging. However, it remains an open question whether the SAT procedure can be strictly improved. This naturally leads to the concept of admissibility, which provides a formal criterion for optimality. We first give the definition of admissibility below.
Definition 1.
A function $f$ is called a level-$\alpha$ deterministic set merging function if it takes a collection of $L$ uncertainty sets along with their miscoverage guarantee $\{(\mathcal{C}_{1},\alpha_{1}),\ldots,(\mathcal{C}_{L},\alpha_{L})\}$ and outputs a single set $\mathcal{C}_{\alpha}$ that satisfies $\mathbb{P}(Y\notin\mathcal{C}_{\alpha})\leq\alpha$, without introducing any external randomness. A level-$\alpha$ deterministic set merging function $f$ is called admissible if there does not exist another level-$\alpha$ deterministic set merging function $g\neq f$ such that $g(\cdot)\subseteq f(\cdot)$ for any valid inputs.
The following theorem establishes that the SAT procedure is not just a valid method, but that its structure is necessary for any admissible approach under general dependence.
Theorem 2.
Under general dependence, every admissible level-$\alpha$ deterministic set merging function can be represented in the form of SAT (Algorithm 1), with synthetic e-value and a convex combination in the aggregation step.
Remark 2.
The converse of Theorem 2 is not true. In Remark C.1 of the Supplement we give a counterexample where SAT with synthetic e-values and certain choice of convex combination is not admissible.
The main idea in proving Theorem 2 is to exploit the duality between uncertainty sets and the e-function defined in (2). This duality allows us to recast the problem of merging uncertainty sets as that of merging e-functions. This perspective enables us to draw on established results from the e-value literature for deeper insight. A key step in the proof of Theorem 2 is to show that all admissible e-function mergers that map collections of e-functions to a single e-function must take the form of convex combinations. We emphasize that this result is not a direct consequence of Theorem 1 in Wang, (2025), as our focus here is on merging functions of a specific form rather than general random variables. In fact, our proof of Theorem 2 uses different techniques from those in the existing e-value literature.
In the special case where all $L$ initial uncertainty sets share the same miscoverage level, it is natural to consider symmetric set merging functions—those that are invariant under permutations of the input sets. The following theorem establishes the admissibility of SAT under this setting. It implies that the merged set produced by any valid and symmetric deterministic function must contain the set generated by the SAT procedure with an arithmetic mean aggregator.
Theorem 3.
When the initial sets are known to be independent and $Y$ is a fixed parameter, the characterization of admissible procedures becomes more intricate. Unlike the dependent case where a single class of functions (convex combinations) is admissible, here the optimal function’s structure depends on the specific values of $(\alpha_{1},\ldots,\alpha_{L})$ and $\alpha$.
The core of the challenge is the discrete nature of the problem. When constructing an admissible function, a key step is to find a function $F:\prod_{\ell=1}^{L}\{0,1/\alpha_{\ell}\}\to\{0,1/\alpha\}$ that “exhausts the error budget”, i.e., $\mathbb{E}\{F(\boldsymbol{e}^{\dagger})\}=1$, where $\boldsymbol{e}^{\dagger}$ is the random vector with independent components $e^{\dagger}_{\ell}$ taking value $1/\alpha_{\ell}$ with probability $\alpha_{\ell}$ and 0 otherwise. The output of $F$ can be written as $F(\boldsymbol{e})=\alpha^{-1}\operatorname{\mathds{1}}\{\boldsymbol{e}\in A\}$ for some set $A\subseteq\prod_{\ell=1}^{L}\{0,1/\alpha_{\ell}\}$. The condition $\mathbb{E}\{F(\boldsymbol{e}^{\dagger})\}=1$ is thus equivalent to $\mathbb{P}(\boldsymbol{e}^{\dagger}\in A)=\alpha$. However, the set of attainable probabilities for this event is a finite, discrete set determined by sums of products of the $\alpha_{\ell}$. If the target level $\alpha$ does not belong to this set, no such function $F$ exists, and a simple sufficient condition for admissibility cannot be applied.
Despite this complexity, a complete characterization of admissibility for the independent case is possible. An admissible merging function corresponds to an optimal rule for partitioning the $2^{L}$ possible outcomes of the synthetic e-values. A merging function is admissible if the set $A$ cannot be enlarged any further without violating the procedure’s validity. This leads to a clear condition: a function is inadmissible if there is an outcome currently not in $A$ but can be moved to $A$ while the function remains valid.
While this provides a theoretical criterion, it confirms that the optimal function depends on the entire discrete probability distribution defined by the $\alpha_{\ell}$ and $\alpha$ values. The full technical details and a formal statement are provided in Section F of the Supplement. This nuanced theoretical landscape explains why there is no single, uniformly optimal aggregator for the independent case. Despite this ambiguity, our simulation studies suggest that the SAT procedure with the average of pairwise products ($k=2$ in (3)) provides an effective and powerful approach in practice.
3 Randomized Set Merging with Synthetic p-values
3.1 Synthetic p-values
While the deterministic approach using e-values provides strong admissibility guarantees, randomized procedures can offer advantages in certain settings, often yielding smaller merged sets. A more traditional test statistic that naturally accommodates randomization is the p-value. A p-value $p\in[0,1]$ is a random variable that satisfies $\mathbb{P}(p\leq t)\leq t$ for all $t\in(0,1)$ under the null hypothesis. As with e-values, p-values can be inverted to form uncertainty sets. In our data-light setting, however, the true “oracle p-value” used to generate the initial set is unavailable. We therefore construct a randomized synthetic p-value that mimics the behaviour of the oracle. For each uncertainty set $\mathcal{C}_{\ell}$ with control level $\alpha_{\ell}$, we define the synthetic p-value as